Position of a particle moving along x-axis is given by x = 3t - 4t2 + t3, where x is in meters and t in seconds.

 

(a)Find the position of the particle at t = 2 s.

 

(b)Find the displacement of the particle in the time interval from t = 0 to t = 4 s.

 

(c)Find the average velocity of the particle in the time interval from t = 2s to t=4s.

 

(d)Find the velocity of the particle at t = 2 s.

 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Pankaj
 
 

(a) x(t) = 3t - 4t2 + t3

=> x(2) = 3 x 2 - 4 x (2)2 + (2)3 = 6 - 4 x 4 + 8 = -2m.

 (b) x(o) = 0

X(4) = 3 x 4 - 4 x (4)2 + (4)3 = 12 m.

Displacement = x(4) - x(0) = 12 m.

(c) = X(4)X(2)/(4-2) = (12-(-2))/2 m/s = 7 m/s

(d) dx/dt = 3 - 8t + 3t2

v(2) (dx/dt)2 = 3 - 8 x 2 + 3 x (2)2 = -1m/s